OpenStudy (vannah):

can anyone help me 3/2(x+2)=x-4

OpenStudy (anonymous):

is it 3/2 *(x+2) or 3/(2(x+2))...? do you divide by (x+2) or multiply?

OpenStudy (anonymous):

if multiply... then: 3x/2 +3/2 *2 = x-4 3x/2 +3 -x=-4 x/2 =-7 x=-14